OpenStudy (anonymous):

Find the solution of this system of linear equations. Separate the x- and y-values with a comma. 4x + 8y =83 8x + 7y = 76 -8x – 16y = -166 8x + 7y = 76 It said it was the adding method I think..

OpenStudy (anonymous):

ellimination method

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x + 8y =83 8x + 7y = 76 find a way where, you will end-up elliminating x

OpenStudy (anonymous):

I posted the wrong question..

OpenStudy (anonymous):

Solve this system of linear equations. Separate the x- and y-values with a comma. -7x = -94 -8y 5x = -70 – 8y this is the question

syed98 (syedmohammed98):

5x = -70 + or - 8y?